OpenStudy (anonymous):

Divison Fractions 5 Grade

OpenStudy (anonymous):

Which problem is solved using this model?

OpenStudy (anonymous):

@jeffyblood

OpenStudy (jeffyblood):

are there any choices?

OpenStudy (anonymous):

4

OpenStudy (jeffyblood):

can you give me the choices?

OpenStudy (anonymous):

ya hold on

OpenStudy (jeffyblood):

k

OpenStudy (anonymous):

\[A.15\div \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

\[B.\frac{ 1 }{ 5 }\div15\]

OpenStudy (moneybaby980):

A

OpenStudy (anonymous):

\[C.3\div \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Really? A

OpenStudy (anonymous):

I have not even said all anwsers

OpenStudy (anonymous):

ok I have 4 Qustions left Im going to open another thing

OpenStudy (moneybaby980):

ok